OpenStudy (nickymail48):

y=-3x 4x+y=2 solve by graphing solve for y

OpenStudy (anonymous):

https://www.desmos.com/calculator

OpenStudy (nickymail48):

ok do u know how to solve without calculator?

OpenStudy (anonymous):

y = -3x y = -4x +2 Put x = 0 , 1 , 2 for few points to graph y.

OpenStudy (michele_laino):

please, note that 1) if we set x=0 in your equation, we get y=2, and so the point (0,2) 2) if we set y=0 in your equation, we get x=1/2, and the point (1/2,0) so the graph of your equation: 4x+y=2, is a line joining the points A=(0,2) and B=(1/2,0) each other

OpenStudy (michele_laino):

sorry an infinite line, of course

OpenStudy (michele_laino):

the graph of the equation y=-3x, is an infinite line which passes through the origin of the coordinate system and the point (1,-3)